For the reaction A → B + C when [A] is plotted versus the time in minutes a straight line is obtained whose slope is -0.011 M/min. What is the concentration of A (in M) after 11.00 min if [A]o = 0.80 M?
Answer #1

if we plot:

[A] vs t

then

this must be ZERO order

in zero order reactions

Zero = C vs t;

For zero order, there is no dependency of concentrations:

dC/dt = k*C^0

dC/dt = k

When developed:

C = C0 + kt

if x axis is "time" then the slope is "k", and y-intercept is initial concentration C0. y-axis if C (concentration)

then

slope = k

C0 = 0.80 M

t = 11 min

substitute

C = C0 + kt

C = 0.80+ (-0.011)(11)

C = 0.679 M